# Comparison of Offset and Ratio Weighted Regressions in Tweedie Models with Application to Mid-Term Cancellations

### v1

@@ 4.4 Observed Gaps Comparison
where $\widehat{\zeta}^{R}_{i}$ and $\widehat{\zeta}^{O}_{i}$ represent the estimated of $\zeta$ from the ratio and offset approaches, respectively. These values, $\widehat{\zeta}^{R}_{i}$ and $\widehat{\zeta}^{O}_{i}$ can also be viewed as realizations of the premium estimators in the ratio and offset approaches.
Our objective is to demonstrate that the observed gaps are equal to zero in the ratio approach when applied to a homogeneous portfolio. However, under the same conditions, alignment between aggregate loss costs and the sum of estimated premiums does not occur in the offset approach. For a heterogeneous portfolio, the observed gaps in both the offset and ratio approaches are not exactly zero due to the influence of the log-link function in premium modeling. Nonetheless, the relationship between the sum of estimated premiums and the aggregate loss costs in the ratio approach can be expressed mathematically. This formulation offers valuable insights into the characteristics and magnitude of the observed gap in this approach.
Our strategy for examining these gaps involves analyzing the MLE for each approach under both homogeneous and heterogeneous portfolio across contracts. By maintaining the assumption of $n$ independent contracts and applying consistent notation and assumptions from previous sections, we ensure coherence across models and approaches.
@@ 4.4.1 Homogeneous Portfolio
Thus, the MLE of $\zeta$ , denoted by $\widehat{\zeta}$, is obtained as follows: $\widehat{\zeta}=\frac{\sum_{i=1}^{n}w_{i}z_{i}}{\sum_{i=1}^{n}w_{i}}$.
@@
    Offset Approach: In the offset approach, with $Z_{i}$ as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}^{2-p}$. Consequently, the MLE for $\zeta$, denoted by $\widehat{\zeta}^{O}$, is:
@@
    The sum of estimated premiums is then calculated as:
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}$ $=\sum_{i=1}^{n}t_{i}\left(\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}\right)$
    $=\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}\sum_{i=1}^{n}t_{i}.$
    We observe that the estimated premiums sum matches the aggregate observed loss cost only under the conditions that $t_{i}=1$ or $p=1$. Given our assumption that the variance parameter $21<p<$ 2 and $t_{i}<1$, it follows that:
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}\neq\sum_{i=1}^{n}t_{i}z_{i}.$
@@
    Ratio Approach: In the ratio approach, where $Z_{i}$ also serves as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}$. The MLE for $\zeta$, denoted by $\widehat{\zeta}^{R}$, is calculated as:
@@
    This result suggests that the ratio approach inherently ensures that the sum of estimated premiums aligns with the aggregate observed loss costs, making it advantageous for achieving balance at the portfolio level.
@@ 4.4.2 Heterogeneous Portfolio
We now compare the approaches in the case of a heterogeneous portfolio with explanatory variables. We then incorporate the characteristics vector $\mathbf{X}_{i}$ to model the premium for each contract $i$, allowing for contract-specific premium estimation and accounting for variation across contracts. Thus, the premium in this approach is represented by Equation (4.1). Unlike the homogeneous portfolio, where a closed-form solution for the MLE of $\beta$ was feasible, here the MLE for the parameter vector $\beta$ does not have a closed-form solution due to its dependence on $\mathbf{X}_{i}$. Consequently, we apply the IRLS algorithm (see Section 2.4.1) to obtain the MLE of $\beta$ for both the offset and ratio approaches, as discussed in previous sections.
To better understand the gap between the aggregate observed loss costs and the sum of estimated premiums in each approach, we initialize the IRLS algorithm with values of $\beta$ based on the homogeneous portfolio assumption. This initialization promotes consistency in comparing the offset and ratio approaches and clarifies how incorporating $\mathbf{X}_{i}$ affects the alignment of estimated premiums with aggregate loss costs.
@@
    $\beta_{0}=\log\left(\widehat{\zeta}^{O}\right),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \beta_{j}=0,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ j=1,\ldots,q.$
    By denoting the initial value by $\beta^{O}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\beta^{O}_{(K)}$, is obtained as follows:
    $\beta^{O}_{(K)}=K\beta^{O}_{(0)}+\underbrace{\sum_{k=1}^{K-1}\left(I^{O}_{n}\left(\beta^{O}_{(k)}\right)\right)^{-1}\nabla^{O}\left(\beta^{O}_{(k)}\right)}_{\delta^{O}_{K-1}}=K\beta^{O}_{(0)}+\delta^{O}_{K-1},$
@@
    $\widehat{\zeta}^{O}_{i(K)}=\left(\widehat{\zeta}^{O}\right)^{K}\exp\left\{\mathbf{X}^{T}_{i}\delta^{O}_{K-1}\right\},$
@@
    $\beta_{0}=\log\left(\widehat{\zeta}^{R}\right),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \beta_{j}=0,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ j=1,\ldots,q.$
    By denoting the initial value by $\beta^{R}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\beta^{R}_{(K)}$, is obtained as follows:
    $\beta^{R}_{(K)}=K\beta^{R}_{(0)}+\underbrace{\sum_{k=1}^{K-1}\left(I^{R}_{n}\left(\beta^{R}_{(k)}\right)\right)^{-1}\nabla^{R}\left(\beta^{R}_{(k)}\right)}_{\delta^{R}_{K-1}}=K\beta^{R}_{(0)}+\delta^{R}_{K-1},$
@@
    $\widehat{\zeta}^{R}_{i(K)}=\left(\widehat{\zeta}^{R}\right)^{K}\exp\left\{\mathbf{X}^{T}_{i}\delta^{R}_{K-1}\right\},$
@@ 4.4.3 Numerical Illustration
To further compare the observed gaps in both regression approaches, especially for the heterogeneous portfolio where an explicit analytical solution is not available, we conduct a numerical example. The following assumptions are made:
@@
    Portfolio Setup: In the homogeneous case, we consider $n=100$ independent contracts with no covariates. In the heterogeneous case, we include two risk factors in the mean parameter of each contract, generated from binomial distributions:
    • $\mathbf{x_{1}}\sim\text{Binomial}(100,0.75)$
    • $\mathbf{x_{2}}\sim\text{Binomial}(100,0.15)$
@@
    Risk Exposure: Risk exposures are randomly drawn from a uniform distribution over the interval $\left[\frac{30}{365},\frac{335}{365}\right]$, and then sorted in increasing order to assign each contract a corresponding rank $i$.
@@
    Loss Cost Scenarios:
    • Increasing Scenario: Loss costs increase with risk exposure: $y_{i}=i$, for $i=1,\ldots,n$.
    • Decreasing Scenario: Loss costs decrease with risk exposure: $y_{i}=n-i-1$, for $i=1,\ldots,n$.
This small numerical example allows us to highlight some of the analytical results discussed earlier. More importantly, it helps to better understand the behavior of the observed gaps for a heterogeneous portfolio, which may not always be expressed analytically based on the approaches (as discussed in Section 4.4.2). Thus, for the Increasing and the Decreasing scenarios, we define the individual gap for contract $i$ using the following equations:
$\Delta_{i}^{R}=t_{i}\left(z_{i}-\widehat{\zeta}^{R}_{i}\right)\quad\text{and}\quad\Delta_{i}^{O}=t_{i}\left(z_{i}-\widehat{\zeta}^{O}_{i}\right).$
Both $\Delta_{i}^{R}$ and $\Delta_{i}^{O}$ were already used in equation (4.14). For risk exposure in $\left[\frac{30}{365},\frac{335}{365}\right]$, the differences $\Delta_{i}$ are plotted as shown in Figure 2. Depending on the portfolio (homogeneous or heterogeneous), the approach (offset or ratio), and the scenario (increasing or decreasing) being studied, many conclusions can be made from Figure 2.
  • Ratio Approach (with solid and dashed blue lines)
    We begin the analysis by recalling that, as shown in Section 4.4.1, $\sum_{i=1}^{n}\Delta_{i}^{R}=0$ for the homogeneous portfolio. This corresponds to the blue dashed line in both graphs of Figure 2. In comparison, the two curves (solid blue lines) representing the heterogeneous portfolio case are very close to the homogeneous case, indicating that $\sum_{i=1}^{n}\Delta_{i}^{R}$ for the heterogeneous case is expected to be near zero. However, the small jumps observed in both solid blue lines suggest that full financial equilibrium cannot be fully achieved.
  • Offset Approach (with solid and dashed red lines)
    We showed in Section 4.4.1 that the offset approach does not allow for $\sum_{i=1}^{n}\Delta_{i}^{O}=0$, which is exactly what we observe with all the red lines of the homogeneous portfolio. Since the sum of the values of the dashed blue line equals zero, the upward (or downward) shift of the red line indicates that financial equilibrium is not achieved. We also observe the same small jumps for the solid blue line associated with the heterogeneous portfolio. In the scenario where costs increase with exposure, it is interesting to note that the offset approach line is shifted upwards compared to the ratio approach line, while the opposite shift is observed in the decreasing scenario. Thus, we can see that the gap in financial equilibrium for the offset approach will be either positive or negative, depending on whether we are in an increasing or decreasing cost-exposure scenario.
Figure 2: Observed gap by risk exposure (left: Scenario 1, right: Scenario 2)
@@ 5.1.3 Covariates Description
The dataset contains various characteristics for each vehicle and contract. To examine the impact of segmentation on pricing, we selected nine key characteristics as covariates, denoted as $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$, $X_{4}$, $X_{5}$, $X_{6}$, $X_{7}$, and $X_{8}$ for confidentiality purposes. $X_{0}$ represents a covariate with a constant value of 1 across all contracts, while the other covariates ($X_{1}$ through $X_{8}$) are binary variables. These covariates capture conventional risk factors such as the insured’s gender, age, vehicle usage, and type of vehicle, in line with standard actuarial practices. To simplify the model, we excluded risk factors that are not typically used in pricing models. As such, we believe the resulting pricing model reflects standard industry practices. Descriptive statistics for these covariates are provided in the appendix.

### v2

@@ 4.3 Sum of Individual Observed Gaps Comparison
where $\widehat{\zeta}_{i}^{O}$ and $\widehat{\zeta}_{i}^{R}$ are the estimated premiums obtained from the offset and ratio approaches. These quantities can be interpreted as realizations of the corresponding premium estimators. Our quantities of interest are therefore the aggregated gaps
$\sum_{i=1}^{n}\Delta_{i}^{O}\quad\text{and}\quad\sum_{i=1}^{n}\Delta_{i}^{R}.$ (4.3)
Although these sums are generally not equal to zero—reflecting the use of a logarithmic link function rather than the canonical link of the Tweedie model in premium modeling—we show that, under the ratio approach, the sum of individual observed gaps tends to be closer to zero than under the offset approach.
To analyze this criterion, we rely on the same simulation framework as in the previous section and examine, for each replication with $p=1.5$, the ratio between the total predicted premiums and the total observed losses under both approaches. This comparison is illustrated in Figure 2. Across all considered sample sizes, the distribution associated with the ratio approach is markedly more concentrated around the value 1—corresponding to a situation close to financial balance—than the distribution obtained under the offset approach. In contrast, the offset approach exhibits substantially greater dispersion, reflecting higher variability around the global balance condition, even as the sample size increases.
A complementary perspective consists in identifying, for each simulation, which approach yields a total predicted premium closer to the observed total losses. According to this criterion, and for all scenarios under consideration, the ratio approach is closer to financial balance in approximately 87% of the replications. This result further supports the view that, from the standpoint of aggregate financial balance, the ratio approach displays a more stable and robust behavior than the offset approach in finite samples.
Figure 2: Financial balance density for n=1,000n=1{,}000, 5,0005{,}000, 20,00020{,}000, and 50,00050{,}000 under the offset and ratio approaches.
Nevertheless, rather than relying solely on simulation results, we argue that additional insight into this financial balance criterion can be gained by examining the parameter estimation equations themselves. To this end, we first consider a homogeneous setting in which the sum of individual observed gaps is exactly zero under the ratio approach. We then extend the analysis to a heterogeneous setting in order to compare the behavior of this aggregate gap under the ratio and offset approaches.
@@ 4.3.1 Homogeneous Portfolio
Thus, the quasi-maximum likelihood estimator (QMLE) of $\zeta$, denoted by $\widehat{\zeta}$, is obtained as follows: $\widehat{\zeta}=\frac{\sum_{i=1}^{n}w_{i}z_{i}}{\sum_{i=1}^{n}w_{i}}$.
@@
    Offset Approach: Under the offset approach, the QMLE of the homogeneous risk parameter $\zeta$ is given by
@@
    The corresponding aggregate estimated premium is therefore
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}=\left(\sum_{i=1}^{n}t_{i}\right)\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}.$ (4.7)
    Exact financial balance, i.e.
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}=\sum_{i=1}^{n}t_{i}z_{i},$
    holds if and only if the $t_{i}^{2-p}$-weighted average of the observed loss costs $z_{i}$ coincides with their $t_{i}$-weighted average. This condition is automatically satisfied in degenerate cases such as $p=1$, or when all exposures $t_{i}$ are equal, in which case the two weighting schemes coincide. However, under the general setting considered in this paper, where $1<p<2$ and exposures satisfy $t_{i}\leq 1$ with heterogeneous durations, there is no structural mechanism in the offset model that enforces this identity. As a result, aggregate financial balance is not guaranteed under the offset approach and may fail in general, although it can still occur for particular realizations of the portfolio. Consequently, the aggregate deviation
    $\sum_{i=1}^{n}\Delta_{i}^{O}=\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}-\sum_{i=1}^{n}t_{i}z_{i}$
    is not constrained to vanish.
@@
    Ratio Approach: In the ratio approach, where $Z_{i}$ also serves as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}$. The QMLE for $\zeta$, denoted by $\widehat{\zeta}^{R}$, is calculated as:
@@
    So, we conclude that $\sum_{i=1}^{n}\Delta_{i}^{R}=0$. This result suggests that the ratio approach inherently ensures that the sum of estimated premiums aligns with the aggregate observed loss costs, making it advantageous for achieving balance at the portfolio level.
@@ 4.3.2 Heterogeneous Portfolio
The homogeneous framework can be extended to account for heterogeneity by incorporating the covariate vector $\mathbf{X}_{i}$ in the premium model for each contract $i$. This allows for contract-specific premium estimation and captures variation across contracts. In contrast to the homogeneous portfolio, where a closed-form expression for the QMLE of $\bm{\beta}$ was available, the presence of covariates implies that the QMLE of the parameter vector $\bm{\beta}$ no longer admits a closed-form solution. This is due to its dependence on $\mathbf{X}_{i}$. Consequently, the QMLE of $\bm{\beta}$ is obtained using the IRLS algorithm (see Section 2.4.1) for both the offset and ratio approaches, as discussed in the previous sections.
To better understand the gap between the aggregate observed loss costs and the sum of estimated premiums in each approach, we initialize the IRLS algorithm with values of $\bm{\beta}$ based on the homogeneous portfolio assumption. This initialization promotes consistency in comparing the offset and ratio approaches and clarifies how incorporating $\mathbf{X}_{i}$ affects the alignment of estimated premiums with aggregate loss costs.
@@
    β_0 = log(^ζ^O), and β_j = 0, for j=1,…,q.
    By denoting the initial value by $\bm{\beta}^{O}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\bm{\beta}^{O}_{(K)}$, is obtained as follows:
    β^O_(K) = K β^O_(0) + ⏟∑_k=1^K - 1 (I^O_n(β^O_(k)))^-1 ∇^O(β^O_(k) )_δ^O_K - 1 = K β^O_(0) + δ^O_K - 1,
@@
    ^ζ^O_i(K) = (^ζ^O)^K exp{ X^T_i δ^O_K - 1},
@@
    β_0 = log(^ζ^R), and β_j = 0, for j=1,…,q.
    By denoting the initial value by $\bm{\beta}^{R}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\bm{\beta}^{R}_{(K)}$, is obtained as follows:
    β^R_(K) = K β^R_(0) + ⏟∑_k=1^K - 1 (I^R_n(β^R_(k)))^-1 ∇^R(β^R_(k) )_δ^R_K - 1 = K β^R_(0) + δ^R_K - 1,
@@
    ^ζ^R_i(K) = ( ^ζ^R)^K exp{ X^T_i δ^R_K - 1},
@@ 4.3.3 Numerical Illustration
To better understand the implications of the parameter estimation equations underlying the two approaches—particularly in the heterogeneous case, where no explicit analytical solution is available—we propose a simple numerical illustration. The objective of this illustration is not to assess the statistical performance of the estimators, but rather to highlight the structural differences between the estimation equations and to clarify how these differences translate into distinct behaviors with respect to the financial balance criterion.
The setup is as follows:
@@
    Portfolio setup. In the homogeneous case, we consider a portfolio of $n=100$ independent contracts without covariates. In the heterogeneous case, two risk factors are incorporated into the mean specification of each contract. These covariates are generated independently from binomial distributions:
    • $x_{1}\sim\text{Binomial}(100,0.75)$,
    • $x_{2}\sim\text{Binomial}(100,0.15)$.
    This specification induces heterogeneity in the portfolio while remaining sufficiently simple to allow a transparent interpretation of the estimation equations.
@@
    Risk exposure. Exposure levels $t_{i}$ are drawn from a uniform distribution on the interval
    $\left[\tfrac{30}{365},\tfrac{335}{365}\right],$
    and then sorted in increasing order to assign contract indices $i=1,\ldots,n$. This ordering reflects a common actuarial situation in which higher indexed contracts correspond to larger exposure levels.
@@
    Loss cost. Rather than assuming a stochastic loss-generating mechanism, we consider a deterministic specification given by
    $y_{i}=i,\qquad i=1,\ldots,n.$
    This choice is deliberate and aims to isolate the effect of the estimation equations themselves. By eliminating random noise, the resulting comparison focuses exclusively on the structural differences between the offset and ratio approaches and on their implications for aggregate financial balance.
Figure 3: Individual observed gaps: single heterogeneous realization (left) and across multiple heterogeneous realizations (right).
We then apply the corresponding parameter estimation equations to this portfolio. For a single realization, the individual observed gaps are displayed in the left panel of Figure 3. This panel allows a direct comparison across portfolio types (homogeneous versus heterogeneous) and estimation approaches (offset versus ratio). Several qualitative insights can be drawn from this figure:
  • Ratio approach (solid and dashed blue curves).
    Recall that, for a homogeneous portfolio, the ratio approach satisfies the exact aggregate balance condition $\sum_{i=1}^{n}\Delta_{i}^{R}=0$, as established in Section 4.3.1. Although this aggregate quantity is not explicitly reported in the figure, the dashed blue curve corresponds to the individual observed gaps obtained under the homogeneous assumption and therefore sums to zero by construction.
    This homogeneous benchmark provides a natural reference for interpreting the solid blue curve, which represents the individual observed gaps obtained under the heterogeneous specification. The two curves are remarkably close over the entire range of exposure levels, indicating that the introduction of covariates does not materially alter the aggregate balance property of the ratio approach. In particular, this visual proximity suggests that the sum $\sum_{i=1}^{n}\Delta_{i}^{R}$ in the heterogeneous case remains close to zero.
    At the same time, the small local deviations and step-like variations observed along the solid blue curve reflect the fact that exact financial balance cannot generally be achieved in the heterogeneous setting, except in accidental cases where the sum of these local variations happens to cancel out. These discrepancies arise from the contract-specific nature of the estimated premiums and from the iterative nature of the estimation procedure. Nevertheless, the overall behavior remains strongly anchored to the homogeneous equilibrium, highlighting the robustness of the ratio approach with respect to portfolio heterogeneity.
  • Offset approach (solid and dashed red curves).
    As established in Section 4.3.1, the offset approach does not satisfy an exact aggregate balance condition, that is, $\sum_{i=1}^{n}\Delta_{i}^{O}\neq 0.$ This feature is directly visible in Figure 3, where both red curves display a systematic displacement relative to the ratio approach, which serves as a natural benchmark since its aggregate gap is zero by construction.
    In the homogeneous case (dashed red curve), this displacement is already apparent, as the offset curve lies uniformly above its ratio counterpart, indicating a departure from financial equilibrium. In the heterogeneous case (solid red curve), the offset curve remains centered around its homogeneous counterpart, suggesting that the imbalance persists even in the presence of additional local variability arising from contract-specific covariates.
To further assess the impact of parameter estimation on financial balance, we consider a Monte Carlo experiment based on 1,000 heterogeneous portfolios generated from varying realizations of $x_{1}$ and $x_{2}$. The corresponding results are reported in the right panel of Figure 3. In this panel, all 1,000 individual gap curves are displayed simultaneously as semi-transparent lines, providing a direct visualization of the outcomes associated with heterogeneous portfolios. To summarize this variability, a pointwise 95% envelope is added in the form of a shaded ribbon, while the bold solid curve represents the median gap across replications. We can see that the individual gap curves obtained across replications are strongly concentrated around their respective homogeneous benchmarks for both approaches. This is reflected by the close proximity of the median heterogeneous curve to the corresponding homogeneous curve shown in the left panel, as well as by the relatively narrow width of the 95% ribbons over the entire range of risk exposure. These patterns indicate that heterogeneity primarily introduces dispersion around the homogeneous structure, rather than inducing systematic shifts in the gap curves themselves.
An important implication of this concentration phenomenon is that the aggregate imbalance observed under the homogeneous specification provides a meaningful first-order diagnostic for the heterogeneous case. Since heterogeneity mainly adds random fluctuations around the homogeneous benchmark, a substantial lack of financial balance at the homogeneous level is unlikely to be systematically corrected once heterogeneity is introduced. Conversely, when the homogeneous imbalance is small, near-balance under heterogeneous modeling becomes more plausible, albeit still not structurally enforced. In this respect, the two approaches differ fundamentally. Exact aggregate balance holds by construction for the homogeneous ratio approach, whereas no such property holds for the homogeneous offset approach. As a result, the concentration of heterogeneous gap curves around their homogeneous counterparts implies that, in heterogeneous portfolios, the ratio approach is typically closer to financial equilibrium than the offset approach.
Finally, we note that the offset approach may occasionally exhibit a better financial balance than the ratio approach for specific realizations, as observed in the simulation results reported in Figure 2, where it is selected in approximately 13% of the replications. As illustrated in Figure 3, such outcomes can partly be attributed to more pronounced random fluctuations, whose aggregate effect may incidentally compensate for the residual imbalance. However, these situations are also more likely to occur in scenarios where the homogeneous offset specification itself happens to be close to financial balance. In such cases, the additional dispersion introduced by heterogeneity may suffice to produce a near-balanced aggregate outcome under the offset approach. Since this compensation mechanism is not structural but contingent on the particular data realization and on the initial homogeneous imbalance, the ratio approach generally provides a more reliable and robust financial balance across heterogeneous portfolios.
@@ 5.1.3 Covariates Description
The dataset contains various characteristics for each vehicle and contract. To examine the impact of segmentation on pricing, we selected nine key characteristics as covariates, denoted as $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$, $X_{4}$, $X_{5}$, $X_{6}$, $X_{7}$, and $X_{8}$ for confidentiality purposes. $X_{0}$ represents a covariate with a constant value of 1 across all contracts, while the other covariates ($X_{1}$ through $X_{8}$) are binary variables. These covariates capture conventional risk factors such as the insured’s gender, age, vehicle usage, and type of vehicle, in line with standard actuarial practices. To simplify the model, we excluded risk factors that are not typically used in pricing models. As such, we believe the resulting pricing model reflects standard industry practices. Descriptive statistics for these covariates are provided in Figure 6.